find the area of the largest circle which can be cut from a square of 4edge 4 in. what is the area of the material wasted?
Answer by mananth(16946)

square edge = 4 in.
Area of square = L^2
= 4*4
= 16 sq.in
...
the center of the square will be the center of the circle.
so radius = 2
..
area of circle = pi*r^2
A= pi*2^2
A=4pi
A= 12.56 sq.in
..
The wasted area = area of square - area of circle
=16-12.56
=3.44 sq.in
...
m.ananth@hotmail.ca